OpenStudy (anonymous):

solve for y 9y+33 =0

OpenStudy (anonymous):

y = -33/9

OpenStudy (anonymous):

can u show work please?

OpenStudy (anonymous):

Sure. I put the constant (33) on the other side by subtracting it so i could get the y term by itsself. If i subtract 33 on one side, i must do it to the other side so that the equation stays equal. so now we have 9y = -33. Then we divide by 9 on both sides to remove the coefficient of y (9) and we get 7 = -33/9

OpenStudy (anonymous):

*y = -33/9 my bad typo

OpenStudy (anonymous):

thanks broski